Fused quartz possesses an exceptionally low coefficient of linear expansion, 5.50 × 10-7 (C°)–1. Imagine you had a bar of fused quartz of length 3.45 m at 20.0°C. By how much, in millimeters, would the bar expand if you heated it to 237°C?Fused quartz possesses an exceptionally low coefficient of linear expansion, 5.50 x 10% (Co)-Imagine you had a bar of fused quartz flength 3.45 m at 20.0℃. By how much, in millimeters, would the bar expand if you heated it to 237°C? Number
Answer #1

(sto) ( t6) = (L-3.us ) (3.ys 237-20) (sSo)(面7)(148.65) L-3.ys =
Answer #2

The length change ΔL from given initial length L0 due to temperature change ΔT is

ΔL=αL0ΔT


ΔT=TT0

Substitute this in the expression for ΔL to obtain the requested change of length in terms of given quantities.

ΔL=αL0(TT0)

For the numerical answer, insert the data and convert the result to millimeters, as requested.


answered by: Muhammad Aslam